Jskajs:

After 3 years , Remington earned $390 in simple interest from a CD into which he Initially deposited $4000. What was the annual interest rate of the CD?

misty34:

@Ultrilliam This question is in the wrong subject

Ferredoxin4:

Using the given information, you can formulate a representative equation. 4000(r)^3=4390 I chose 4390 instead of 390 as it's c umulative Now solve for r. 4000(r)^3=4390 r^3=1.0975 r=1.0315 The interest rate is about 3.15%